# TempoPFN: Synthetic Pre-training of Linear RNNs for Zero-shot Time Series Forecasting

### v2

@@ Abstract
Foundation models for zero-shot time series forecasting face challenges in efficient long-horizon prediction and reproducibility, with existing synthetic-only approaches underperforming on challenging benchmarks. This paper presents TempoPFN, a univariate time series foundation model based on linear Recurrent Neural Networks (RNNs) pre-trained exclusively on synthetic data. The model uses a GatedDeltaProduct architecture with state-weaving for fully parallelizable training across sequence lengths, eliminating the need for windowing or summarization techniques while maintaining robust temporal state-tracking. Our comprehensive synthetic data pipeline unifies diverse generators—including stochastic differential equations, Gaussian processes, and audio synthesis—with novel augmentations. In zero-shot evaluations on the Gift-Eval benchmark, TempoPFN achieves top-tier competitive performance, outperforming all existing synthetic-only approaches and surpassing the majority of models trained on real-world data, while being more efficient than existing baselines by leveraging fully parallelizable training and inference. We open-source our complete data generation pipeline and training code, providing a reproducible foundation for future research.
@@ 1 Introduction
Recent advances in large language models have inspired foundation models for time series forecasting that enable zero-shot predictions across diverse datasets without fine-tuning (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024; Auer et al., 2025). By treating historical observations as input context, these models democratize forecasting for non-experts and excel in data-scarce domains.
However, current approaches face critical limitations. Transformer-based models struggle with long-horizon forecasting due to quadratic complexity and error accumulation (Zeng et al., 2023). While non-linear RNNs like those in TiReX (Auer et al., 2025) maintain temporal state, they require sequential processing that limits scalability. Although some recent models attempt synthetic-only pre-training including ForecastPFN (Dooley et al., 2023), CauKer (Xie et al., 2024), and Mamba4Cast (Bhethanabhotla & Swelam, 2024) none reported state-of-the-art performance on the Gift-Eval benchmark. TabPFN-TS (Hoo et al., 2024), which adapts a tabular foundation model to time series, achieves strong Gift-Eval performance but does not release its synthetic pre-training data, limiting reproducibility and extensibility.
We introduce TempoPFN (see Table 1 and Figure 1), a time series forecasting foundation model using linear RNNs with GatedDeltaProduct recurrence (Siems et al., 2025) for parallelizable training and inference across the sequence length. Unlike TiRex (Auer et al., 2025) which argued that non-linear RNNs like sLSTM are necessary for time-series forecasting due to their state-tracking capabilities we find that linear RNNs based on the GatedDeltaProduct recurrence are sufficient, in line with recent research demonstrating how linear RNNs can perform state-tracking (Grazzi et al., 2025). Our synthetic data pipeline unifies diverse generators with novel augmentations, ensuring exclusive synthetic pre-training to prevent benchmark leakage. Unlike TabPFN-TS, we open-source our complete data generation pipeline and training code as a basis for future research (available at https://github.com/automl/TempoPFN).
@@
  • The TempoPFN architecture, to our knowledge, the first univariate time series foundation model based on linear RNNs with GatedDeltaProduct recurrence. Our architecture and input representation allows the prediction of all future time-stamps in parallel, producing coherent quantile forecasts, without patching or windowing heuristics. We further propose a state-weaving mechanism for linear RNNs that facilitates bidirectional information flow across horizons without overhead.
  • We design a synthetic data pipeline combining existing and novel synthetic generators with a cascade of augmentations, ensuring diverse temporal structures without relying on real-world data, thereby eliminating benchmark leakage and mitigating privacy concerns associated with training on real-world data. We release a fully open-source synthetic data generation pipeline for time series forecasting that achieves competitive competitive performance on Gift-Eval.
  • Compared to nonlinear RNNs and transformer time series foundation models, TempoPFN achieves top-tier competitive zero-shot performance on Gift-Eval, surpassing all other synthetic-only approaches and the vast majority of models trained on real-world data. This result is achieved without any non-linearity in the recurrence, demonstrating that linear RNNs as a scalable and powerful alternative to non-linear RNNs and transformers for time series foundation models.
@@ 2 Background and Related Work
Time Series Forecasting. Time series forecasting aims to predict future values $y_{T+1:T+H}$ from historical observations $y_{1:T}$. Traditional methods such as ARIMA (Box & Jenkins, 1968) and exponential smoothing (Hyndman et al., 2008) typically produce point estimates, while probabilistic forecasting captures uncertainty by modeling the predictive distribution $p(y_{T+1:T+h}\mid y_{1:T})$. The advent of deep learning has significantly expanded this toolkit, introducing transformers (Vaswani et al., 2017) and modern recurrent architectures (Beck et al., 2024; Gu & Dao, 2023).
A major recent development is zero-shot forecasting, where models pre-trained on diverse corpora can directly predict unseen time series without requiring fine-tuning. This paradigm mirrors the transformative shift seen in natural language processing and computer vision, where foundation models enable efficient cross-domain adaptation without costly per-task training.
Most successful zero-shot approaches leverage transformer architectures. Chronos (Ansari et al., 2024), TimesFM (Das et al., 2024), and MOIRAI (Woo et al., 2024) use techniques such as patching, frequency-specific projections, and masked modeling to handle heterogeneous time series data. Building on this foundation, MOIRAI-MOE (Liu et al., 2025) incorporates a sparse mixture-of-experts architecture to achieve token-level specialization and improved robustness. Among true zero-shot models, MOIRAI currently demonstrates state-of-the-art performance on Gift-Eval while carefully avoiding overlap with evaluation benchmarks.
An alternative approach focuses on synthetic data pretraining. ForecastPFN (Dooley et al., 2023) trains exclusively on synthetic distributions featuring multi-scale trends, seasonality, and Weibull noise, enabling Bayesian zero-shot inference through single forward passes. TimePFN (Taga et al., 2025) extends this framework to multivariate scenarios using Gaussian process kernels and linear coregionalization. Similarly, TabPFN-TS (Hoo et al., 2024) represents time series in tabular format and leverages TabPFNv2 (Hollmann et al., 2025), achieving competitive performance despite the limited availability of its underlying synthetic data.
@@
Recent work has also revisited recurrent architectures for long-horizon forecasting. TiRex (Auer et al., 2025), currently the top performer on Gift-Eval, uses xLSTM (Beck et al., 2024) pre-trained on synthetic Gaussian processes, Chronos datasets, and carefully selected Gift-Eval subsets, enhanced with data augmentation techniques including amplitude modulation, censoring, and spike injection. In contrast, TempoPFN takes a different approach by exploiting linear RNNs with GatedDeltaProduct mechanisms and negative eigenvalues, enabling fully parallelizable training without requiring patching or summarization while relying exclusively on synthetic pretraining data to eliminate real-world leakage concerns.
Linear RNNs.
Recurrent neural networks have seen a resurgence in interest with the emergence of linear RNNs also known as state-space models. While non-linear RNNs are non-trivially parallelizable (Gonzalez et al., 2024), linear RNNs can be parallelized using a chunk-wise parallel form (Yang et al., 2024a) or an associative scan (Gu & Dao, 2023; Martin & Cundy, 2018). Formally, linear RNN layers transform input sequences ${\bm{x}}_{1:t}\in\mathbb{R}^{l}$ into outputs $\hat{{\bm{y}}}_{1:t}\in\mathbb{R}^{p}$ through the recurrence
${\bm{H}}_{i}={\bm{A}}({\bm{x}}_{i}){\bm{H}}_{i-1}+{\bm{B}}({\bm{x}}_{i})\text{~with output~}\hat{{\bm{y}}}_{i}=\mathrm{dec}({\bm{H}}_{i},{\bm{x}}_{i})\text{ for~}i\in\{1,\ldots,t\}$ (1)
@@
where ${\bm{A}}:\mathbb{R}^{l}\to\mathbb{R}^{n\times n}$ parameterizes the state-transition matrix, ${\bm{B}}:\mathbb{R}^{l}\to\mathbb{R}^{n\times d}$ governs state inputs, and $\mathrm{dec}:\mathbb{R}^{n\times d}\times\mathbb{R}^{l}\to\mathbb{R}^{p}$ produces the layer output. Variants such as Mamba (Dao & Gu, 2024), GLA (Yang et al., 2024a), and mLSTM (Beck et al., 2024) adopt diagonal transitions, while others explore richer parameterizations. More expressive formulations relax diagonal state-transition constraints, as seen in DeltaNet (Schlag et al., 2021; Irie et al., 2023; Yang et al., 2024b), TTT-Linear (Sun et al., 2024), RWKV-7 (Peng et al., 2025), B’MOJO (Zancato et al., 2024), and Titans (Behrouz et al., 2024).
Within this framework, we use DeltaProduct (Siems et al., 2025) as our token-mixing mechanism, which generalizes DeltaNet’s non-diagonal transitions by expressing ${\bm{A}}({\bm{x}}_{i})$ as a product of $n_{h}$ generalized Householder transformations, enabling a rank-$n_{h}$ transformation of the matrix-valued hidden state ${\bm{A}}({\bm{x}}_{i})=\prod_{j=1}^{n_{h}}\bigl({\bm{I}}-\beta_{i,j}\,{\bm{k}}_{i,j}{\bm{k}}_{i,j}^{\top}\bigr)$. For each token ${\bm{x}}_{i}$, the model generates $n_{h}$ normalized keys ${\bm{k}}_{i,j}=\psi({\bm{W}}_{j}{\bm{x}}_{i})/\|\psi({\bm{W}}_{j}{\bm{x}}_{i})\|_{2}$, values ${\bm{v}}_{i,j}={\bm{V}}_{j}{\bm{x}}_{i}$, and coefficients $\beta_{i,j}=\phi({\bm{U}}_{j}{\bm{x}}_{i})$ using learnable matrices ${\bm{W}}_{j},{\bm{V}}_{j},{\bm{U}}_{j}$, SiLU activation $\psi$ (Hendrycks & Gimpel, 2016), and a sigmoid-based gating function $\phi$. Siems et al. (2025) found that increasing $n_{h}$ leads to significantly improved length-extrapolation, language modeling, and state-tracking on permutation tasks, all capabilties which are equally desirable for time-series forecasting.
@@ 3.1 Architecture
Figure 2: The TempoPFN architecture (3 blocks), using stacked GatedDeltaProduct blocks, learnable initial states ${\bm{H}}_{0}^{i}$ and state-weaving.
@@
Input representation. TempoPFN uses an input representation in which history (timesteps + values) and future (timesteps) are concatenated into one token sequence enabling communication between future time-steps for coherent predictions. In contrast to TiReX, which presummarizes time-steps into windows of size 32, TempoPFN directly operates on the individual time-steps. Each time step $t_{i}$ is encoded using GluonTS (Alexandrov et al., 2020) time features (e.g., seasonality indicators, day-of-week, or index-based encodings) that are linearly projected into the embedding dimension of the model. For historical steps, observed values $y_{i}$ are projected via a linear layer, while missing values are handled by a learnable NaN embedding. The historical embedding is obtained by additively combining the value embedding and the time-feature embedding. For future time steps, only the time-feature embedding is used.
Backbone. The core of TempoPFN is a stack of 10 encoder layers, each based on the Gated DeltaProduct block from the flash-linear-attention library (Yang & Zhang, 2024), originally derived from the LLaMA architecture (Touvron et al., 2023). Each block consists of three components: (1) token mixing through a Gated DeltaProduct recurrence with short one-dimensional convolutions (kernel size 16–32), (2) pre-normalization applied before the recurrent unit to stabilize training, and (3) a gated MLP for channel-wise feature transformation. This design combines the parallelization advantages of linear recurrences with the expressivity of lightweight convolutional and feedforward operations. Non-causality via state weaving. Whereas DeltaProduct was originally developed for autoregressive language modeling, forecasting across a full horizon does not require causal masking. To exploit this property, we introduce state weaving. Specifically, the final hidden state of each layer ${\bm{H}}_{t}^{i}$ is added to the learnable initial state of the next layer ${\bm{H}}_{0}^{i+1}$. This mechanism enables bidirectional information flow across the entire sequence length without additional parameters or computational overhead through explicit bidirectionality (Hwang et al., 2024; Afzal et al., 2025), effectively lifting the causal constraint while maintaining computational efficiency.
@@
Prediction. At the output stage, embeddings corresponding to the forecast horizon are extracted from the final encoder block. These embeddings are passed through a linear projection head that outputs multiple quantiles of the predictive distribution, enabling probabilistic forecasting. Overall, this design allows to directly predict all future values given a history using a single forward pass.
@@ 3.2 Synthetic Data Generation
To train our time series foundation model, we generated a large and diverse dataset using 10 different synthetic generators. This approach combines established data generation techniques with novel methods to capture a wide spectrum of temporal patterns and behaviors. For a more comprehensive description of each generator refer to Appendix A.
@@
Existing Generators. We adapted several established generators from prior work to ensure comprehensive coverage of common temporal patterns. The ForecastPFN generator (Dooley et al., 2023) composes multiplicative trend and seasonality components, combining linear and exponential growth terms with sinusoidal harmonics. The generator includes Weibull-distributed noise and augmentations such as time warping, magnitude scaling, and spike injection, with filtering mechanisms to avoid extreme values. KernelSynth, following Ansari et al. (2024), samples univariate time series from Gaussian process priors with composite kernels. Base kernels include periodic (ExpSineSquared), stationary (RBF, RationalQuadratic), and noise (WhiteKernel) components, combined through addition or multiplication to yield smooth yet varied trajectories. We extended this approach with a broader Gaussian Process generator, as in Bhethanabhotla & Swelam (2024) that randomly combines kernels with greater functional diversity, producing wider ranges of stationary and nonstationary patterns. The CauKer generator (Xie et al., 2024) introduces causal dependencies by sampling from structural causal models (SCMs). Each node represents a Gaussian process with composite kernels and stochastic mean functions, while edges in a random DAG apply nonlinear transformations. We generate 21-channel multivariate series, treating each channel as an independent univariate signal to capture diverse, interdependent dynamics.
@@
Novel Generators. We developed several new generators to fill gaps in existing approaches and capture specific temporal behaviors. Sawtooth creates ramp-like patterns with upward or downward slopes, enhanced with small linear trends and low-amplitude seasonal components to avoid overly idealized signals. Step Function produces piecewise constant series with configurable changepoints, step sizes, and drift, using Gaussian smoothing at boundaries along with added noise, seasonality, and anomalies. For anomaly-rich data, we created two specialized generators. Anomaly produces baseline signals with periodic or clustered spikes, varying in magnitude regimes (constant, trending, cyclical, or correlated random) and timing patterns. Spikes emphasizes event-driven behavior by placing sharp spikes on flat baselines, with configurable shapes (V, inverted V, or plateau variants) arranged in bursty or evenly spread patterns. The Sine Wave generator provides clean oscillatory patterns with configurable period, amplitude, phase, and noise, offering fundamental periodic signals for learning basic oscillatory structures. To capture highly complex, real-world dynamics, we introduce Audio-Inspired Generators that use procedural audio synthesis techniques implemented with Pyo (Belanger, 2016). These generators model phenomena such as Stochastic Rhythms for event data, Financial Volatility with market shocks and clustering, Network Topology with traffic bursts and congestion, and Multi-Scale Fractals for self-similar patterns. Our most sophisticated contribution is the stochastic differential equation (SDE) generator, a flexible synthetic data generator based on a regime-switching, time-inhomogeneous Ornstein–Uhlenbeck (OU) process. The OU process follows the SDE $dy_{t}=\theta(t,r_{t})\,\bigl(\mu(t,r_{t})-y_{t}\bigr)\,dt+\sigma(t,r_{t})\,dW_{t}$ where $\theta(t,r_{t})$ is the mean reversion speed, $\mu(t,r_{t})$ the time-varying mean, and $\sigma(t,r_{t})$ the volatility. Each parameter depends on both time $t$ and a latent regime $r_{t}\in\{0,1\}$ that evolves as a Markov chain. This framework enables parameters to shift abruptly across regimes while drifting smoothly over time through polynomial, sinusoidal, logistic, or piecewise-linear trends. Seasonal patterns are injected additively into both mean and volatility components, with amplitudes subject to gradual growth or decay. For enhanced realism, we optionally replace standard Brownian motion with fractional Brownian motion, introducing long-memory dynamics through the Hurst exponent $H\in(0,1)$. Each simulated series undergoes global rescaling and shifting before additive Gaussian measurement noise is applied. This construction produces highly diverse temporal structures, capturing regime shifts, non-stationarity, periodicity, and measurement noise within a principled stochastic framework.

### v3

@@ Abstract
Foundation models for zero-shot time series forecasting face challenges in efficient long-horizon prediction and reproducibility, with existing synthetic-only approaches underperforming on challenging benchmarks. This paper presents TempoPFN, a univariate time series foundation model based on linear Recurrent Neural Networks (RNNs) pre-trained exclusively on synthetic data. The model uses a GatedDeltaProduct architecture with state-weaving for fully parallelizable training across sequence lengths, eliminating the need for windowing or summarization techniques while maintaining robust temporal state-tracking. Our comprehensive synthetic data pipeline unifies diverse generators, including stochastic differential equations, Gaussian processes, and audio synthesis, with novel augmentations. In zero-shot evaluations on the Gift-Eval, fev-bench and Chronos-ZS benchmarks, TempoPFN achieves top-tier competitive performance, outperforming all existing synthetic-only approaches and surpassing the majority of models trained on real-world data, while being more efficient than existing baselines by leveraging fully parallelizable training and inference. We open-source our complete data generation pipeline and training code, providing a reproducible foundation for future research.
@@ 1 Introduction
Recent advances in large language models have inspired foundation models for time series forecasting that enable zero-shot predictions across diverse datasets without fine-tuning (ansari2024chronos; das2024decoder; woo2024unified; auer2025tirex). By treating historical observations as input context, these models democratize forecasting for non-experts and excel in data-scarce domains.
However, current approaches face critical limitations. Transformer-based models struggle with long-horizon forecasting due to quadratic complexity and error accumulation (zeng2023transformers). While non-linear RNNs like those in TiReX (auer2025tirex) maintain temporal state, they require sequential processing that limits scalability. Although some recent models attempt synthetic-only pre-training including ForecastPFN (dooley2023forecastpfn), CauKer (cauker), and Mamba4Cast (mamba4cast) none reported state-of-the-art performance on the Gift-Eval benchmark. TabPFN-TS (hoo2024the), which adapts a tabular foundation model to time series, achieves strong Gift-Eval performance but does not release its synthetic pre-training data, limiting reproducibility and extensibility.
We introduce TempoPFN (see Table 1 and Figure 1), a time series forecasting foundation model using linear RNNs with GatedDeltaProduct recurrence (siems2025deltaproduct) for parallelizable training and inference across the sequence length. We adopt the Prior-Data Fitted Network (PFN) framework (muller2022transformers), treating zero-shot forecasting as Bayesian inference approximated via in-context learning on a diverse synthetic prior (see Appendix A.1 for a detailed background). Unlike TiRex (auer2025tirex) which argued that non-linear RNNs like sLSTM are necessary for time-series forecasting due to their state-tracking capabilities, we find that linear RNNs based on the GatedDeltaProduct recurrence are sufficient, in line with recent research demonstrating how linear RNNs can perform state-tracking (grazzi-iclr25a). As detailed in Appendix B.1, DeltaProduct applies orthogonal rotations via multiple online gradient steps, enabling superior state-tracking compared to diagonal SSMs. Our synthetic data pipeline unifies diverse generators with novel augmentations, ensuring exclusive synthetic pre-training to prevent benchmark leakage. Unlike TabPFN-TS, we open-source our complete data generation pipeline and training code as a basis for future research (available at https://github.com/automl/TempoPFN).
@@
  • The TempoPFN architecture is, to our knowledge, the first univariate time series foundation model based on linear RNNs with GatedDeltaProduct recurrence. Our architecture and input representation allows the prediction of all future time-stamps in parallel, producing coherent quantile forecasts, without patching or windowing heuristics. We further propose a state-weaving mechanism for linear RNNs that facilitates bidirectional information flow across horizons without overhead.
  • We design a synthetic data pipeline combining existing and novel synthetic generators with a cascade of augmentations, ensuring diverse temporal structures without relying on real-world data, thereby eliminating benchmark leakage and mitigating privacy concerns associated with training on real-world data. We release a fully open-source synthetic data generation pipeline for time series forecasting that achieves competitive performance on Gift-Eval.
  • Compared to nonlinear RNNs and transformer time series foundation models, TempoPFN achieves top-tier competitive zero-shot performance on Gift-Eval, fev-bench and Chronos-ZS, surpassing all other synthetic-only approaches and the vast majority of models trained on real-world data. This result is achieved without any non-linearity in the recurrence, demonstrating that linear RNNs as a scalable and powerful alternative to non-linear RNNs and transformers for time series foundation models.
@@ 2 Background and Related Work
Time Series Forecasting. Time series forecasting aims to predict future values $y_{T+1:T+H}$ from historical observations $y_{1:T}$. Classical methods such as ARIMA (box1968some) and exponential smoothing (hyndman2008forecasting) produce point estimates, while probabilistic forecasting models the predictive distribution $p(y_{T+1:T+h}\mid y_{1:T})$. Deep learning has expanded this toolkit with transformers (vaswani2017attention) and modern recurrent architectures (beck2024xlstm; gu2023mamba). A key recent development is zero-shot forecasting, in which models pre-trained on diverse corpora can predict unseen time series without fine-tuning, mirroring cross-domain generalization in NLP and vision. Most successful approaches employ transformers: Chronos (ansari2024chronos), TimesFM (das2024decoder), and MOIRAI (woo2024unified) use patching, frequency-specific projections, and masked modeling to handle heterogeneous data. MOIRAI-MOE (liu2024moirai) adds a sparse mixture of experts for token-level specialization and robustness. Among true zero-shot models, MOIRAI currently achieves state-of-the-art performance on Gift-Eval while avoiding benchmark overlap.
Prior-data Fitted Networks (PFNs) and Synthetic Data. PFNs (muller2022transformers) represent a paradigm shift from solving a single task to learning a universal inference algorithm by training a neural network to approximate the posterior predictive distribution induced by a prior over datasets. By minimizing the expected negative log-likelihood across datasets sampled from this prior (see Appendix A.1 for more details), PFNs enable in-context learning as fast approximate Bayesian inference: instead of storing solutions to tasks, the model’s weights encode an inference algorithm that conditions on a small context (e.g., a time-series history) and predicts future values in a single forward pass (muller2022transformers; hollmann2023tabpfn). As PFN performance is determined by the expressiveness of the prior, models are typically trained on large synthetic corpora. For instance, TabPFN (hollmann2023tabpfn) uses structural causal models for tabular data, ForecastPFN (dooley2023forecastpfn) employs trend–seasonality priors for time series, TimePFN (taga2025timepfn) extends this to multivariate settings with Gaussian process kernels, and TabPFN-TS (hoo2024the) adapts PFNs to time series via TabPFNv2 (hollmann2025accurate). In this paper, the prior is our synthetic temporal-dynamics pipeline (see Section 3.2). Training on this synthetic prior enables TempoPFN to perform zero-shot probabilistic forecasting on unseen time series.
@@
Linear RNNs and State-Space Models. Recent work has revisited recurrent architectures for long-horizon forecasting. TiRex (auer2025tirex) uses xLSTM (beck2024xlstm) pre-trained on synthetic Gaussian processes, Chronos datasets, and selected Gift-Eval subsets, combined with augmentations such as amplitude modulation, censoring, and spike injection. In contrast, TempoPFN leverages linear RNNs with GatedDeltaProduct (siems2025deltaproduct) mechanisms and negative eigenvalues (grazzi-iclr25a), enabling fully parallelizable training without patching or summarization, while relying solely on synthetic pretraining data to avoid any leakage. Linear RNNs have regained interest due to their efficient parallelization. While non-linear RNNs are not easily parallelizable (NEURIPS2024_0b2b199f), linear RNNs admit chunk-wise parallelization (yang-icml24a) or associative scans (gu2023mamba; martin2018parallelizing). Formally, they map input sequences ${\bm{x}}_{1:t}\!\in\!\mathbb{R}^{l}$ to outputs $\hat{{\bm{y}}}_{1:t}\!\in\!\mathbb{R}^{p}$ via
${\bm{H}}_{i}={\bm{A}}({\bm{x}}_{i}){\bm{H}}_{i-1}+{\bm{B}}({\bm{x}}_{i}),\qquad\hat{{\bm{y}}}_{i}=\mathrm{dec}({\bm{H}}_{i},{\bm{x}}_{i}),\quad i=1,\ldots,t,$ (1)
@@
where $\bm{A}$} parameterizes state transitions, $\bm{B}$} state inputs, and $\mathrm{dec}$ the output. Variants such as Mamba (dao-icml24a), GLA (yang-icml24a), and mLSTM (beck2024xlstm) use diagonal transitions, while more expressive models relax this constraint, including DeltaNet (schlag-icml21a; irie2023practical; yang-neurips24a), TTT-Linear (sun-arxiv24a), RWKV-7 (peng2025rwkv7gooseexpressivedynamic), B’MOJO (zancato-neurips24a), and Titans (behrouz2024titans).
@@ 3.1 Architecture
Figure 2: The TempoPFN architecture (3 blocks) uses stacked GatedDeltaProduct blocks, learnable initial states ${\bm{H}}_{0}^{i}$ and state-weaving.
@@
Input representation. TempoPFN uses an input representation in which history (time steps + values) and future (time steps) are concatenated into one token sequence enabling communication between future time steps for coherent predictions. In contrast to TiReX, which presummarizes time steps into windows of size 32, TempoPFN directly operates on the individual time steps. Each time step $t_{i}$ is encoded using GluonTS (gluonts) time features (e.g., seasonality indicators, day-of-week, or index-based encodings) that are linearly projected into the embedding dimension of the model. For historical steps, observed values $y_{i}$ are projected via a linear layer, while missing values are handled by a learnable NaN embedding. The historical embedding is obtained by additively combining the value and the time-feature embeddings. For future time steps, only the time-feature embedding is used.
Backbone. The core of TempoPFN is a stack of 10 encoder layers, each based on the Gated DeltaProduct block (siems2025deltaproduct) from the flash-linear-attention library (fla), originally derived from the LLaMA architecture (touvron2023llama). Each block consists of three components: (1) token mixing through a Gated DeltaProduct recurrence with short one-dimensional convolutions (kernel size 16–32), (2) pre-normalization applied before the recurrent unit to stabilize training, and (3) a gated MLP for channel-wise feature transformation. This design combines the parallelization advantages of linear recurrences with the expressivity of lightweight convolutional and feedforward operations. DeltaProduct, which generalizes DeltaNet’s non-diagonal transitions by expressing ${\bm{A}}({\bm{x}}_{i})$ in Equation 1 as a product of $n_{h}$ generalized Householder transformations, enabling a rank-$n_{h}$ transformation of the matrix-valued hidden state ${\bm{A}}({\bm{x}}_{i})=\prod_{j=1}^{n_{h}}\bigl({\bm{I}}-\beta_{i,j}\,{\bm{k}}_{i,j}{\bm{k}}_{i,j}^{\top}\bigr)$. For each token ${\bm{x}}_{i}$, the model generates $n_{h}$ normalized keys ${\bm{k}}_{i,j}=\psi({\bm{W}}_{j}{\bm{x}}_{i})/\|\psi({\bm{W}}_{j}{\bm{x}}_{i})\|_{2}$, values ${\bm{v}}_{i,j}={\bm{V}}_{j}{\bm{x}}_{i}$, and coefficients $\beta_{i,j}=\phi({\bm{U}}_{j}{\bm{x}}_{i})$ using learnable matrices ${\bm{W}}_{j},{\bm{V}}_{j},{\bm{U}}_{j}$, SiLU activation $\psi$ (hendrycks2016gaussian), and a sigmoid-based gating function $\phi$. siems2025deltaproduct found that increasing $n_{h}$ leads to significantly improved length-extrapolation, language modeling, and state-tracking on permutation tasks, all capabilities which are equally desirable for time-series forecasting.
Non-causality via state weaving. Whereas DeltaProduct was originally developed for autoregressive language modeling, forecasting across a full horizon does not require causal masking. To exploit this property, we introduce state weaving. Specifically, the final hidden state of each layer ${\bm{H}}_{t}^{i}$ is added to the learnable initial state of the next layer ${\bm{H}}_{0}^{i+1}$. This mechanism enables bidirectional information flow across the entire sequence length without additional parameters or computational overhead through explicit bidirectionality (NEURIPS2024_c7f795dc; afzal2025linear), hence allowing future time-steps to attend to the entire history and future context, preventing the information bottleneck typical of causal RNNs during the prediction phase.
@@
Prediction. At the output stage, embeddings corresponding to the forecast horizon are extracted from the final encoder block. These embeddings are passed through a linear projection head that outputs multiple quantiles of the predictive distribution, enabling probabilistic forecasting (see also Appendix D for details on the quantile loss used for training). Overall, this design allows to directly predict all future values given a history using a single forward pass.
@@ 3.2 Synthetic Data Generation
To train our time series foundation model, we generated a large and diverse dataset using 10 different synthetic generators. This approach combines established data generation techniques with novel methods to capture a wide spectrum of temporal patterns and behaviors. For a more comprehensive description of each generator, refer to Appendix C.
@@
Existing Generators. We adapted several established generators from prior work to ensure comprehensive coverage of common temporal patterns. The ForecastPFN generator (dooley2023forecastpfn) composes multiplicative trend and seasonality components, combining linear and exponential growth terms with sinusoidal harmonics. The generator includes Weibull-distributed noise and augmentations such as time warping, magnitude scaling, and spike injection, with filtering mechanisms to avoid extreme values. KernelSynth, following ansari2024chronos, samples univariate time series from Gaussian process priors with composite kernels. Base kernels include periodic (ExpSineSquared), stationary (RBF, RationalQuadratic), and noise (WhiteKernel) components, combined through addition or multiplication to yield smooth yet varied trajectories. We extended this approach with a broader Gaussian Process generator, as in mamba4cast that randomly combines kernels with greater functional diversity, producing wider ranges of stationary and nonstationary patterns. The CauKer generator (cauker) introduces causal dependencies by sampling from structural causal models (SCMs). Each node represents a Gaussian process with composite kernels and stochastic mean functions, while edges in a random DAG apply nonlinear transformations. We generate 21-channel multivariate series, treating each channel as an independent univariate signal to capture diverse, interdependent dynamics.
@@
Novel Generators. We developed several new generators to fill gaps in existing approaches and capture specific temporal behaviors. Sawtooth creates ramp-like patterns with upward or downward slopes, enhanced with small linear trends and low-amplitude seasonal components to avoid overly idealized signals. Step Function produces piecewise constant series with configurable changepoints, step sizes, and drift, using Gaussian smoothing at boundaries along with added noise, seasonality, and anomalies. For anomaly-rich data, we created two specialized generators. Anomaly produces baseline signals with periodic or clustered spikes, varying in magnitude regimes (constant, trending, cyclical, or correlated random) and timing patterns. Spikes emphasizes event-driven behavior by placing sharp spikes on flat baselines, with configurable shapes (V, inverted V, or plateau variants) arranged in bursty or evenly spread patterns. The Sine Wave generator provides clean oscillatory patterns with configurable period, amplitude, phase, and noise, offering fundamental periodic signals for learning basic oscillatory structures. To capture highly complex, real-world dynamics, we introduce Audio-Inspired Generators that use procedural audio synthesis techniques implemented with Pyo (pyo). These generators model phenomena such as Stochastic Rhythms for event data, Financial Volatility with market shocks and clustering, Network Topology with traffic bursts and congestion, and Multi-Scale Fractals for self-similar patterns. Our most sophisticated contribution is the stochastic differential equation (SDE) generator, a flexible synthetic data generator based on a regime-switching, time-inhomogeneous Ornstein–Uhlenbeck (OU) process. The OU process follows the SDE $dy_{t}=\theta(t,r_{t})\,\bigl(\mu(t,r_{t})-y_{t}\bigr)\,dt+\sigma(t,r_{t})\,dW_{t}$ where $\theta(t,r_{t})$ is the mean reversion speed, $\mu(t,r_{t})$ the time-varying mean, and $\sigma(t,r_{t})$ the volatility. Each parameter depends on both time $t$ and a latent regime $r_{t}\in\{0,1\}$ that evolves as a Markov chain. This framework enables parameters to shift abruptly across regimes while drifting smoothly over time through polynomial, sinusoidal, logistic, or piecewise-linear trends. Seasonal patterns are injected additively into both mean and volatility components, with amplitudes subject to gradual growth or decay. For enhanced realism, we optionally replace standard Brownian motion with fractional Brownian motion, introducing long-memory dynamics through the Hurst exponent $H\in(0,1)$. Each simulated series undergoes global rescaling and shifting before additive Gaussian measurement noise is applied. This construction produces highly diverse temporal structures, capturing regime shifts, non-stationarity, periodicity, and measurement noise within a principled stochastic framework.